OpenStudy (anonymous):

Which of the following is the slope between the points (-2, 4) and (2, -2)?

OpenStudy (anonymous):

The formula to find the slope is m = y2-y1 over x2-x1. So you plug in your values like this: -2-4 over 2-(-2) Once you solve for those it will leave you with a fraction, that will be your answer.

OpenStudy (anonymous):

In your case you will have to simplify the fraction when your done, and then that fraction will be your answer.

OpenStudy (anonymous):

i'M WORKING ON IT

OpenStudy (anonymous):

Lol ok..

OpenStudy (anonymous):

it equals -14/1??

OpenStudy (anonymous):

Mm not quite. The formula should look like this: m = \[\frac{ -2-4 }{ 2-(-2) }\]

OpenStudy (anonymous):

still dont get it....

OpenStudy (anonymous):

The values on the top, let's start there. -2-4 = what??

OpenStudy (anonymous):

2-(-2)

OpenStudy (anonymous):

No. Plug -2-4 into your calculator and then tell me what you get

OpenStudy (anonymous):

Ohhh SORRY -6 LOL...

OpenStudy (anonymous):

Lol no prob. That's your new numerator, ok? Now do the same with the bottom values 2-(-2). Then tell me what you get

OpenStudy (anonymous):

is it 4?

OpenStudy (anonymous):

Yep :) Now that's your denominator. What does your fraction look like right now?

OpenStudy (anonymous):

-6/4

OpenStudy (anonymous):

Great. Now can you simplify your fraction anymore than it already is?

OpenStudy (anonymous):

yass to -3/2

OpenStudy (anonymous):

So your slope or m is = -3/2 :)

OpenStudy (anonymous):

Thank YOU so MUCHHH...

OpenStudy (anonymous):

Lol no problem. :) Don't forget to close your question ^.^